OpenStudy (rissaunnie):

how do i write this in exponential form 6^-3 X 6^-8

OpenStudy (owlcoffee):

You might want to use the property applied to exponents: \[A ^{^-m}=\frac{ 1 }{ A^m }\] Where "A" is any nymber or variable, and "m" is any real number.

OpenStudy (owlcoffee):

Also, exponents with the same base are operable, in this case being multiplication: \[A^m \times A^n = A ^{m+n}\] In this case A is 6, and that means you can add the exponents and keeping that 6 there.